OpenStudy (anonymous):

(4g 2 - 9) ÷ (2g - 3)

OpenStudy (ankit042):

(a^2-b^2)=(a+b)(a-b) now use this to solve

OpenStudy (ankit042):

Hint 4g^2-9 = (2g-3)(2g+3) now complete it

OpenStudy (anonymous):

thanks